OpenStudy (anonymous):

Which shows this equation in standard form? 7 – 3(x – y) = 5x + 2 A. 3x – 8y = 5 B. 8x + 3y = 5 C. 8x – 3y = 5 D. 8x – 3y = –5

OpenStudy (anonymous):

Really stumped on this one.

OpenStudy (texaschic101):

standard form is : Ax + By = C 7 - 3(x - y) = 5x + 2 --- distribute through the parenthesis 7 - 3x + 3y = 5x + 2 -- subtract 7 to both sides -3x + 3y = 5x + 2 - 7 -- subtract 5x from both sides -3x + 3y - 5x = 2 - 7 -- combine like terms -8x + 3y = -5 -- multiply by -1 to make x positive 8x - 3y = 5 ==> standard form